OpenStudy (anonymous):

A recipe calls for 1/3 of a cup of milk for 11 cookies. How many cups of milk are needed to make 165 cookies? A. 6 B. 55 C. 5 D. 11 First person who answers get a medal.

OpenStudy (texaschic101):

(1/3) / 11 = x/165 -- (1/3 cup milk to 11 cookies = x cup milk to 165 cookies) cross multiply 11(x) = (1/3)(165) 11x = 55 -- divide both sides by 11 x = 5 It takes 5 cups of milk to make 165 cookies any questions ?

OpenStudy (anonymous):

Thank you

OpenStudy (texaschic101):

anytime :)